# Evaluation of Audio Compression Codecs

### v1

@@
abstract=true, fontsize=10pt, footinclude=false, headinclude=false, headsepline=false, mpinclude=false, numbers=noenddot, paper=letter, parskip=half-, twocolumn=true, twoside=false, [subfloat]aboveskip=0pt, belowskip=0pt, font=scriptsize,sf, format=hang, labelsep=space, margin=1pt, singlelinecheck=0,
Evaluation of Audio Compression Codecs
Thien T. Duong    Jan P. Springer
@@ Abstract
Perceptual quality of audio is the combination of aural accuracy and listener-perceived sound fidelity. It is how humans respond to the accuracy, intelligibility, and fidelity of aural media. Today this fidelity is also heavily influenced by the use of audio compression codecs for storing aural media in digital form. We argue that, when choosing an audio compression codec, users should not only look at compression efficiency but also consider the sonic perceptual quality properties of available audio compression codecs.
We evaluate several commonly used audio compression codecs in terms of compression performance as well as their sonic perceptual quality via codec performance measurements, visualizations, and PEAQ scores. We demonstrate how perceptual quality is affected by digital audio compression techniques, providing insights for users in the process of choosing a digital audio compression scheme.
@@ 1 Introduction
Perceptual quality of audio is the combination of aural accuracy and listener-perceived sound fidelity [Ellis:PRA:1992]. How listeners experience digital audio depends on sound fidelity, i.e. the accuracy of a digital signal representing the reproduction of sound. High-fidelity or hi-fi digital audio ideally should have low levels of noise and distortion in its signal to provide a high perceptual quality.
Our goal is to assess widely used audio compression codecs such as [FLACWebsite] or AAC and MP3 [Brandenburg:MP3AACE:1999] with respect to compression efficiency as well as how the compression techniques in these audio codecs affect the perceptual quality of the signal reconstruction. In addition, we use visual analytics derived techniques to present the perceptual impact of compression techniques via spectrograms, sound field diagrams, and objective perceptual scoring using Perceptual Evaluation of Audio Quality (PEAQ) [Thiede:PEAQITUSOMPAQ:2000].
We argue that in choosing an audio compression codec, users should not only look at the compression efficiency but also consider the sonic perceptual quality of an audio compression codec.
@@ 2 Related Work
Audio codecs use general compression techniques to reduce the size of digital audio file. Traditional audio compression techniques include Huffman coding [Bosi:IDACS:2003], Linear Predictive Coding (LPC) [Liebchen:IMALC:2004], and psychoacoustic-based perceptual encoding [Burg:DSMCAS:2016].
@@
Audio codecs are categorized into lossy or lossless depending on the compression scheme. Lossy audio codecs, such as MP3 and AAC [Brandenburg:MP3AACE:1999] or Vorbis [Moffitt:OVOFA:2001], compress digital audio by removing unimportant data, i.e. parts of the digital signal that can be difficult to perceive by a listener because it is not within the human hearing threshold. Lossless audio codecs such as [FLACWebsite] compress digital audio by only removing redundancies and compressing repeated patterns in the audio signal while maintaining all of the original signal’s data.
Assessing digital audio codecs requires comparing compression efficiency as well as evaluating the resulting perceptual quality. Compression metrics include compression ratio, encoding speed, and decoding speed. Evaluating perceptual quality requires a more complex assessment. Classically, perceptual quality assessments require trained listeners to perform listening tests [Olive:HHTL:2011]. However, listening tests are a rather subjective assessment with inherent listener bias as well as reproducibility problems. We present perceptual-quality assessment using visualizations as well as objective scoring. Tools such as frequency spectra [AudacityWebsite], spectrograms [SpekWebsite], and sound-field diagrams [IzotopeWebsite] show the impact of compression artifacts to the perceptual quality of encoded digital audio. In addition, Perceptual Evaluation of Audio Quality (PEAQ) is used to quantify the perceptual quality by generating a single Objective Difference Grade (ODG) score of encoded digital audio [Thiede:PEAQITUSOMPAQ:2000].
In recent years, experiments with AI-driven audio compression techniques have been conducted. These techniques are based on well-known neural network architectures such as Residual Vector Quantized Generative Adversarial Networks (RVQGAN) [Kumar:HFACIRVQGAN:2023]. We also compare RVQGAN audio compression technique with traditional audio compression techniques in our evaluation.
@@ 3.1 Hardware and Software
All experiments were conducted on a computer workstation featuring an AMD Ryzen 5 5600X CPU, 32GB RAM at 3200 MHz, and a Samsung 970 EVO NVMe SSD storage device. The reference listening setup included a [DangerousMusicWebsite] digital-to-analog converter (DAC), a [SangakuDIYWebsite], and [SennheiserHD600Website] open-back headphones.
Encoding and decoding of digital audio used [FMediaWebsite], a flexible audio processing tool supporting most existing lossy and lossless codecs. Additionally, fmedia also allowed recording of file sizes as well as elapsed time for audio encoding and audio decoding processes. These measurements were used to calculate compression ratio, audio encoding speed, and audio decoding speed.
@@ 3.2 Audio Codecs and Settings
The lower the compression ratio, the more storage space a digital audio file will occupy. Encoding speed and decoding speed are expressed in samples per microsecond and calculated as shown in equation 2.
@@
The most commonly used audio codecs were chosen for our experiments: FLAC, MP3, AAC, and Vorbis. These audio codecs have different settings, allowing users to choose different compression levels. To keep our experiments close to real life scenarios, the most common settings were chosen for each audio codec. Table 1 provides an overview of the settings used for the various audio encoding formats.
@@
Lossless FLAC Compression 0 – 8 6
@@
AAC VBR 1 – 5 5
Vorbis VBR 1 – 10 7
Table 1: Audio-codec settings. The Range column shows an option’s range while the Settings column shows the used setting. CBR stands for constant bit rate, VBR stands for variable bit rate.
@@ 3.3 Visualization Tools
To visualize perceptual quality, a combination of spectral analysis tools were employed. These include built-in spectrum generator from [AudacityWebsite] (cf. figure 1a), spectrogram analysis by [SpekWebsite] (cf. figure 1b), and sound-field visualizations provided by [IzotopeWebsite] (cf. figure 1c).
@@
Figure 1: Visualization of perceptual audio quality. LABEL:sub@fig:setup:visual:spectrum+sample Audacity generated spectrum shows peak loudness within the highest frequency (48 kHz); x-axis shows time, y-axis shows frequency, colored spectrum indicates the loudness in dB. LABEL:sub@fig:setup:visual:spectrogram Spek generated spectrogram shows loudness range (0 dB to -120 dB) of audio signal’s frequency range; x-axis shows time, y-axis shows frequency range, colored spectrum indicates loudness in dB. LABEL:sub@fig:setup:visual:sound+field Sound field visualization generated by Audacity using the Insight 2 plug-in shows width, depth, and height of the sound stage as well as the location of individual sound sources within the sound stage. The half-circle line indicates front of the sound stage while bright dots indicate stereo images. All visualizations based on Dreams by Fleetwood Mac (2001 remaster vinyl quality, 96 kHz sample rate).
Each type of visualization represents a different aspect of aural perceptual quality. Spectrum analysis shows how audio compression techniques affect the loudness of a digital audio file through its frequency response. Spectrograms display time, frequency, and amplitude with a color range for easy visual differentiation. Sound fields reveals how audio compression techniques affect the stereo image and sound stage, i.e. how audio signals are distributed between the left and the right channel of a stereo system and the width, depth and height of an audio signal, respectively.
@@ 3.4 Perceptual Evaluation of Audio Quality
To quantify the impact of audio-compression techniques on digital audio signals, an objective scoring system is required. Perceptual Evaluation of Audio Quality (PEAQ) [Thiede:PEAQITUSOMPAQ:2000] is an ITU-standardized algorithm to assess the perceptual quality of digital audio. PEAQ generates an ODG score (cf. table 2), expressed in terms of a quality gap between a tested audio signal and its original audio signal.
@@
Table 2: Objective Difference Grade scoring range, after [ITU:BS.1387-2:2023]. A generated score from the objective measurement corresponds to a Subjective Difference Grade (left column). Resolution of ODG scores is limited to one decimal.
Advanced PEAQ (APEAQ) [Kabal:EIITUPEAQ:2002] is a more appropriate version of PEAQ for newer audio codecs [Delgado:CWSUP:2020], while basic PEAQ’s total Noise-to-Mask (totalNMR) model output variable (MOV) is well-correlated with subjective listening tests [Dick:ODAQ:2025]. In our study, we used both APEAQ’s ODG and basic PEAQ’s totalNMR MOV as objective perceptual quality evaluation metrics. totalNMR is measured in negative decibels and, ideally, should be as far as possible from 0 dB while at the same time totalNMR is trending close to 0 dB.
To calculate APEAQ’s ODG and PEAQ’s totalNMR we used GstPEAQ [Holters:GSTPEAQ:2015] and [peaqbGitHub], respectively. It is important to note that these open-source projects do not fully conform to [ITU:BS.1387-2:2023]. However, our pre-study experiments showed that both GstPEAQ as well as peaqb perform within reasonably close limits to each other and should not do worse than subjective listening test [Holters:GSTPEAQ:2015].
@@ 3.5 Test Samples
We conducted our experiments on 50 digital audio files, using high-fidelity conversions of CDs, DVDs, vinyls, or cassettes from a wide range of music genres and epochs. For each sample, we collected the file size, the sampling rate, and the original bit rate used for conversion. Each sample was then encoded using our chosen audio codecs at their respective quality settings (cf. table 1). To showcase the impact of audio-compression techniques on the perceptual quality of digital audio, we chose audio files for visualizations that present high levels of perceptual quality as well as a high degree of accuracy in stereo imaging and sound-stage appearance.
@@ 4.1 Audio-Codec Performance
Table 3 provides excerpts from the dataset for selected entries along with the most common quality setting, compression ratio as a percentage, and encoding speed as well as decoding speed in samples per microsecond.
Track Codec Setting Compr. Ratio Enc. Speed Dec. Speed
@@
Table 3: Excerpt from the codec-performance dataset showing audio codec compression as percentage ratio and encoding as well as decoding speeds in samples per microsecond using the audio codec configurations established in table 1.
The full dataset, [FMediaACCDGithubRepo], currently contains 53 entries with each entry also specifying additional metadata (e.g., genre) as well as its file size and the number of samples. This is followed by compression bit rate, the compressed file size, the compression, and the amount of time as well as the speed for encoding and decoding for each of the tested audio codecs.
Table 4 presents the average compression ratio, encoding speed and decoding speed of the dataset.
Codec Setting Comp. Ratio Enc. Speed Dec. Speed
@@
FLAC, the only lossless compression scheme tested, achieves an average encoding speed of 32 samples per microsecond, i.e. 32 million samples per second. MP3 at 128 kbps, a lossy compression scheme, exhibits both the best average compression ratio at 92% as well as the best decoding speed at 149 samples per microsecond, i.e. 149 million samples per second.
@@ 4.2 Perceptual Quality
Figure 2 visualizes the spectra of all tested audio codecs versus the uncompressed digital audio file of the song Iron Man by Black Sabbath.
Figure 2: Spectra for all tested audio codecs (exported from Audacity). X-axis denotes the frequency range up to 24 kHz. The y-axis shows the loudness range from 0 dB to -150 dB. The color-coded graphs refer to uncompressed (black), FLAC level 6 (red), MP3 CBR 320 kbps (yellow), MP3 CBR 128 kbps (green), AAC CBR 256 kbps (orange), AAC VBR level 5 (teal), and Vorbis VBR level 7 (pink). The graphs represent encoded audio signal outputs for the respective audio codec. Note the spectrum for the uncompressed audio signal is completely aligned with the spectrum of FLAC level 6, showing that FLAC encoded audio exhibits the exact same loudness as the uncompressed audio signal across the entire frequency range.
The graph of FLAC level 6 aligns perfectly with the spectrum of uncompressed audio, showing that the loudness of FLAC encoded audio exactly matches the loudness of the uncompressed audio across the entire spectrum. On the other hand, spectra for all audio encoded with lossy compression techniques deviate from the spectrum of the original audio signal. The same pattern of reduction in loudness can be seen for the interval 15 kHz to above 20 kHz. MP3 320 kbps and MP3 128 kbps elevate the loudness of the entire frequency range before cutoff at 18 kHz and 15 kHz, respectively. AAC CBR 256 kbps loudness reduction happens much earlier in the frequency range, around 7 kHz. To our surprise, AAC VBR level 5 loudness closely matches with the spectra for FLAC and the uncompressed audio very well with only a slight loudness reduction at the high-frequency range. Vorbis level 7 also exhibits loudness reduction at high frequencies. However, unlike the other tested lossy compression techniques, it manifests a smooth decline in loudness.
[SpekWebsite] was used to further investigate and visualize the loudness change in encoded audio compared to uncompressed audio. Figure 3 shows that all audio encoded with lossy compression schemes exhibit very quiet loudness in the range -100 dB to -105 dB, marked dark blue in figure 3a, which is further reduced to near silence at -120 dB, marked black in figures 3c to 3g, for high frequencies above 15 kHz.
@@
MP3 320 kbps retains some portion of very quiet loudness until 20 kHz, while AAC VBR level 5 retains some very quiet loudness. Vorbis level 7 loudness reduction is not as flat as other lossy codecs.
Figure 4 shows example sound-field visualizations for stereo images and sound stages capturing the 3:00 – 3:30 min interval of Iron Man by Black Sabbath, when Tony Iommi performs the guitar solo.
@@
Figure 4: Sound-field visualizations for the track Iron Man by Black Sabbath compressed using different audio encoding schemes. LABEL:sub@fig:evaluation:iron+man+sf:a shows the sound-field visualization from 3:00 – 3:30 min for the uncompressed signal and LABEL:sub@fig:evaluation:iron+man+sf:b – LABEL:sub@fig:evaluation:iron+man+sf:g show the sound-field visualization for the same time interval encoded with FLAC, MP3s, AACs, and Vorbis, respectively. The patterns of dots should lean to the right, capturing the stereo image of the electric guitar’s sound in the recording. Compared to the uncompressed audio signal’s stereo image, the FLAC encoded audio signal was able to accurately capture the stereo image while audio signals from other lossy encoders show some inaccuracies.

### v2

@@ Abstract
Perceptual quality of audio is the combination of aural accuracy and listener-perceived sound fidelity. It defines how humans respond to the accuracy, intelligibility, and fidelity of aural media. Today this fidelity is also heavily influenced by the use of audio-compression codecs used for storing aural media in digital form. We argue that, when choosing an audio-compression codec, users should not only look at compression efficiency but also consider the sonic perceptual quality properties of available audio-compression codecs.
We evaluate several audio-compression codecs, traditional as well as AI-based, in terms of compression performance and their sonic perceptual quality via codec-performance measurements, various visualization techniques, and a family of PEAQ scores. We demonstrate how perceptual quality is affected by digital audio-compression techniques, providing insights for a wide range of users in the process of choosing a digital audio-compression scheme.
@@ 1 Introduction
Perceptual quality of audio is the combination of aural accuracy and listener-perceived sound fidelity [undefl]. How listeners experience digital audio depends on sound fidelity, i.e. the accuracy of a digital signal representing the reproduction of sound. High-fidelity or hi-fi digital audio ideally should have low levels of noise and distortion in its signal to provide a high perceptual reproduction quality.
Our goal is to assess widely used audio-compression codecs such as [undefm] or AAC and MP3 [undefc] as well as AI-based audio codecs such as RVQGAN [undefw] with respect to compression efficiency as well as how the compression techniques in these audio codecs affect the perceptual quality of the signal reconstruction. We use visual analytics derived techniques to present the perceptual impact of compression techniques via spectrograms, sound-field diagrams, and objective perceptual scoring based on Perceptual Evaluation of Audio Quality (PEAQ) [undefaf] and its variant implementations [undefu].
While subjective listening tests are considered the gold standard for audio-quality evaluation, our test method is based on objective quality scoring using PEAQ. We believe this will allow for more straightforward replication by non-domain experts as well as allowing additional results to be added into community maintained repositories.
We argue that in choosing an audio-compression codec, users should not only look at the compression efficiency but also consider the sonic perceptual quality of an audio-compression codec in reproducing audio signals.
@@ 2 Related Work
Audio codecs use general compression techniques to reduce the size of digital-audio content. Traditional audio compression techniques include Huffman coding [undefq, undefb], Linear Predictive Coding (LPC) [undefx], and psychoacoustic-based perceptual encoding [undefd].
@@
Audio codecs are categorized into lossy and lossless depending on the employed compression scheme. Lossy audio codecs, such as MP3 and AAC [undefc] or Vorbis [undefy], compress digital audio by removing unimportant data, i.e. parts of the digital signal that can be difficult to perceive by a listener because it is not within the human hearing threshold. Lossless audio codecs such as [undefm] encode digital audio by only removing redundancies and compressing repeated patterns in the audio signal while maintaining all of the original signal’s data.
Assessing digital audio codecs requires comparing compression efficiency as well as evaluating the resulting perceptual quality. Compression metrics include compression ratio, encoding speed, and decoding speed. Evaluating perceptual quality requires more complex assessments.
Classically, perceptual-quality assessments require trained listeners to perform listening tests [undefz], which is considered the gold standard in perceptual-quality evaluation. A subjective listening test such as MUSHRA [undefs] requires trained listeners to be accurate. However, it has been proven that biases exist in MUSHRA [undefaj]. Additionally, recent media coverage of blind listening tests has brought attention to instances where listeners were unable to discriminate between sound transmitted through significantly different types of conductive materials [undeff]. This emphasizes our view that subjective listening tests do not necessarily match human intuition regarding audio quality.
We present our perceptual-quality assessment using various visualization techniques as well as objective scoring to evaluate model-predicted perceptual impact. Tools such as frequency spectra [undefa], spectrograms [undefad], and sound-field diagrams [undeft] allow exposing the impact of compression artifacts on the perceptual quality of compressed digital audio. Additionally, Perceptual Evaluation of Audio Quality (PEAQ) [undefaf] is used to quantify the perceptual quality of audio signals by generating a single Objective Difference Grade (ODG) score of encoded digital audio.
However, PEAQ is not perfect and there are many attempts to improve on the objective evaluation model. Advanced PEAQ [undefu] has been suggested for modern audio codecs, which tend to exhibit better perceptual encoding algorithms. [undefk] show that by using the PEAQ’s model output variable (MOV) totalNMR by itself is already a good evaluation metric in combination with advanced PEAQ’s ODG score. Additionally, the 2f model [undefv] in particular is a good example for objectively approximating subjective MUSHRA scores [undefj].
In recent years, experiments with audio compression techniques based on machine-learning and deep-learning techniques have been conducted. There have been several AI-based compression technique proposed, which include end-to-end approaches [undefaa], lightweight-transformer approaches [undefi], and novel audio codecs such as APCodec [undef]. We chose [undefw]’s Residual Vector Quantized Generative Adversarial Networks (RVQGAN) model because the authors’ claim that their model outperforms competing traditional audio compression algorithms.
We do not intend to replace subjective listener based evaluation with our work. Rather we explore how computational-perceptual models such as PEAQ and its variants in combination with a variety of visualization techniques can facilitate reproducible benchmarking and exploratory evaluation in line with psychoacoustic principles for scalable audio-codec analysis.
@@ 3.1 Hardware and Software
All experiments were conducted on a computer workstation featuring an AMD Ryzen 5 5600X CPU, 32GB RAM at 3200 MHz, and a Samsung 970 EVO NVMe SSD storage device. The reference listening setup included a [undefh] digital-to-analog converter (DAC), a [undefab], and [undefac] open-back headphones.
Encoding and decoding of digital audio used [undefo], a flexible audio processing tool supporting most existing lossy and lossless codecs. Additionally, fmedia also allows recording of resulting file sizes as well as elapsed time for audio encoding and audio decoding processes. These measurements were used to calculate compression ratio, audio encoding speed, and audio decoding speed.
@@ 3.2 Audio Codecs and Settings
The lower the compression ratio the more storage space a digital-audio file will occupy. Encoding speed and decoding speed are expressed in samples per microsecond and calculated as shown in equation 2.
@@
Lossless FLAC Compression 0 – 8 level 6
@@
AAC VBR 1 – 5 level 5
Vorbis VBR 1 – 10 level 7
Table 1: Default audio-codec settings for the audio codecs used in our tests. The Range column shows an option’s range while the Setting column displays the value used in testing. CBR stands for constant bit rate, VBR stands for variable bit rate.
Instead of finding and using optimal parameters for each audio codec, presets were selected to represent frequently used default values. Our study prioritizes ecological validity and reproducibility over audio-codec tweaking when evaluating real-world audio-codec behavior under general use setups. Therefore, rather than describing theoretical audio-codec restrictions, reported results describe real-world deployment conditions. Table 1 provides an overview of the settings used for the various audio encoding formats.
@@ 3.3 Visualization Tools
Figure 1: Perceptual-audio quality visualization. LABEL:sub@fig:setup:visual:spectrum+sample Audacity generated spectrum shows peak loudness within the highest frequency (48 kHz); x-axis shows time, y-axis shows frequency, colored spectrum indicates loudness in dB. LABEL:sub@fig:setup:visual:spectrogram Spek generated spectrogram shows loudness range (0 dB to -120 dB) of an audio signal’s frequency range; x-axis shows time, y-axis shows frequency range, colored spectrum indicates loudness in dB. LABEL:sub@fig:setup:visual:sound+field Sound-field visualization generated by Audacity using the Insight 2 plug-in shows width, depth, and height of the sound stage as well as the location of individual sound sources within that sound stage. The half-circle line indicates sound-stage front while bright dots indicate stereo images. Visualizations based on Dreams by Fleetwood Mac (2001 remaster vinyl quality, sampled at 96 kHz).
To visualize perceptual quality a combination of spectral analysis tools were employed. These include the built-in spectrum generator of [undefa] (cf. figure 1a), spectrogram analysis by [undefad] (cf. figure 1b), and sound-field visualizations provided by [undeft] (cf. figure 1c). Each type of visualization represents a different aspect of aural perceptual quality. Spectrum analysis shows how audio-compression techniques affect the loudness of a digital-audio signal through its frequency response. Spectrograms display time, frequency, and amplitude with a color range for easy visual differentiation. Sound fields reveal how audio-compression techniques affect the stereo image and the sound stage, i.e. how audio signals are distributed between the left and the right channel of a stereo system and the width, depth, and height of an audio signal, respectively.
@@ 3.4 Perceptual Evaluation of Audio Quality
Table 2: Objective Difference Grade (ODG) scoring range, after [undefr]. A generated score from the objective measurement corresponds to a Subjective Difference Grade (left column) qualitatively expressed as Impairment. ODG score resolution is limited to one decimal.
To quantify the impact of audio-compression techniques on digital audio signals an objective scoring system is required. Perceptual Evaluation of Audio Quality (PEAQ) [undefaf] is an ITU-standardized algorithm to assess the perceptual quality of digital audio. PEAQ generates an Objective Difference Grade (ODG) score (cf. table 2), expressed in terms of a quality gap between a tested audio signal and its respective original audio signal.
There are two versions of PEAQ: basic PEAQ (BPEAQ) and advanced PEAQ (APEAQ) [undefu]. Both use the same Fast Fourier Transform (FFT) model to convert time-domain signals into frequency-domain signals for evaluation. APEAQ additionally includes a filter-bank ear model to accurately simulate the human auditory system’s cochlear response. [undefj] report BPEAQ to be less computational intensive than APEAQ, while APEAQ is the more appropriate version of PEAQ for newer audio codecs such as AAC and Vorbis. This is due to these audio codecs’ perceptual models being more complex. BPEAQ’s model output variable (MOV) total noise-to-mask ratio (totalNMR) is well-correlated with subjective listening tests [undefk]. totalNMR is measured in negative decibels and ideally should be as far as possible from 0 dB while at the same time totalNMR is trending close to 0 dB. In our study, we calculate both basic and advanced PEAQ’s ODG scores as well as basic PEAQ’s totalNMR value to showcase the performance of each scoring metric.
Additional models exist, which improve on PEAQ based on past evaluation. Of these the 2f model [undefv] performs significantly better than others [undefag]. The 2f model uses the ADB and the AvgModDiff1 MOVs of the BPEAQ model to calculate a Mean MUSHRA Score Estimate ($\text{MMS}_{\text{est}}$) [undefv] as shown in equation 3.
[…]
The result of equation 3 is then clamped to the range [0…100] to calculate the actual 2f score as shown in equation 4.
$\text{2f score}=\min(100,\max(0,\text{MMS}_{\text{est}}))$ (4)
Past experiments have shown that the 2f model’s average score is the closest equivalent to a subjective listening test result [undefv].
We use McGill University’s PEAQ Matlab model [undefu] to calculate a BPEAQ ODG score and extract the totalNMR MOV. In addition, ADB and AvgModDiff1 MOVs are extracted to calculate a 2f score. To calculate APEAQ’s ODG score, GstPEAQ [undefp] is utilized. We note that GstPEAQ is an open-source project that does not fully conform to [undefr]. However, pre-study experiments have shown that GstPEAQ performs within reasonably close limit, i.e. 0.03 difference in root mean square error, to [undefr] and is therefore expected to perform similar to a subjective listening test [undefp].
@@ 3.5 Test Samples
We conducted our experiments with 50+ digital-audio files, using high-fidelity conversions of CDs, DVDs, vinyls, and cassettes from a wide range of music genres and epochs. For each sample, we collected the file size, the sampling rate, and the original bit rate used for conversion. Each sample was then encoded using our chosen audio codecs at their respective quality settings (cf. table 1). To showcase the impact of audio-compression techniques on the perceptual quality of digital audio, in the following we chose a set of audio examples for visualizations that present high levels of perceptual quality as well as high degrees of accuracy in stereo imaging and sound-stage appearance.
@@ 4.1 Audio-Codec Performance
Track Codec Setting Compression Ratio Encoding Speed Decoding Speed
@@
Table 3: Excerpt from our codec-performance dataset showing audio-codec compression as percentage ratio and encoding as well as decoding speeds in samples per microsecond using the audio-codec configurations established in table 1.
Table 3 provides excerpts from our dataset for selected entries along with the chosen quality setting, compression ratio as a percentage, and encoding speed as well as decoding speed in samples per microsecond. The full dataset, [undefn], contains 50+ entries. Each entry specifies additional metadata (e.g., genre) as well as original file size, number of digital samples, compression bit rate, compressed file size, original encoding algorithm, and encoding time as well as speed for encoding and decoding for each of the tested audio codecs.
Codec Setting Compression Ratio Encoding Speed Decoding Speed
@@
Table 4 presents average values for compression ratio, encoding speed and decoding speed of our dataset. FLAC, the only lossless compression scheme tested, achieves an average encoding speed of 32 samples per microsecond, i.e. 32 million samples per second. MP3 at 128 kbps, a lossy compression scheme, exhibits both the best average compression ratio at 92% as well as the best decoding speed at 149 samples per microsecond, i.e. 149 million samples per second.
@@ 4.2 Perceptual Quality
Figure 2: Spectra for all tested audio codecs (exported from Audacity). X-axis denotes frequency range up to 24 kHz. Y-axis shows loudness range from 0 dB to -150 dB. Color-coded graphs refer to uncompressed (black), FLAC level 6 (red), MP3 CBR 320 kbps (yellow), MP3 CBR 128 kbps (green), AAC CBR 256 kbps (orange), AAC VBR level 5 (teal), and Vorbis VBR level 7 (pink). Graphs represent encoded audio-signal outputs for the respective audio codec. Note the spectrum for the uncompressed audio signal is completely aligned with that of FLAC level 6, meaning both exhibit the exact same loudness across the entire frequency range. Spectra of both MP3, AAC CBR 256 kbps, and Vorbis encoded audio diverge from uncompressed audio spectrum, revealing the impact of lossy compression schemes on the overall loudness.
Figure 2 visualizes the spectra of all tested audio codecs versus the uncompressed digital audio of Iron Man by Black Sabbath. The graph for FLAC level 6 aligns perfectly with the spectrum of the uncompressed audio signal, showing that the loudness of FLAC encoded audio exactly matches the loudness of the uncompressed audio across the entire spectrum. On the other hand, spectra for all audio encoded with lossy compression techniques deviate from the spectrum of the original audio signal. The same pattern of reduction in loudness can be seen for the interval 15 kHz to above 20 kHz. MP3 320 kbps and MP3 128 kbps elevate the loudness of the entire frequency range before cutoff occurs at 18 kHz and 15 kHz, respectively. AAC CBR 256 kbps loudness reduction happens much earlier in the frequency range, around 7 kHz. To our surprise, AAC VBR level 5 loudness closely matches with the spectra for FLAC and uncompressed audio with only a slight loudness reduction at the high-frequency range. Vorbis level 7 also exhibits loudness reduction at high frequencies. However, unlike the other tested lossy compression techniques it manifests a smooth decline in loudness.
@@
[undefad] was used to further investigate and visualize the loudness change in encoded audio compared to uncompressed audio. Figure 3 shows that all audio encoded with lossy compression schemes exhibit very quiet loudness in the range -100 dB to -105 dB, marked dark blue in figure 3a, which is further reduced to near silence at -120 dB, marked black in figures 3c to 3g, for high frequencies above 15 kHz. MP3 at 320 kbps maintains much of the original structure but exhibits slight attenuation in the upper frequencies, while MP3 at 128 kbps shows a more pronounced high-frequency cutoff and reduced spectral detail. AAC and Vorbis demonstrate similar high-frequency constraints and reshape frequency-dependent energy in the mid-to-lower frequency. The redistribution of quantization noise is part of AAC’s and Vorbis’ perceptual encoding strategy to compress audio content while making sure the audio quality is perceptually masked [undefe].
@@
Figure 4: Sound-field visualizations for the track Iron Man by Black Sabbath compressed using different audio encoding schemes. LABEL:sub@fig:evaluation:iron+man+sf:a shows the sound-field visualization from 3:00 – 3:30 min for the uncompressed audio signal and LABEL:sub@fig:evaluation:iron+man+sf:b – LABEL:sub@fig:evaluation:iron+man+sf:g show the sound-field visualization for the same time interval encoded with FLAC, MP3, AAC, and Vorbis, respectively. Dot patterns should lean to the right, capturing the stereo image of the electric guitar’s sound in the recording. Compared to the uncompressed audio signal’s stereo image, FLAC encoded audio was able to accurately capture the stereo image while other lossy encoders exhibit some inaccuracies.